OpenStudy (bluepatriot):

Can somebody help me figure this out and show me the steps so that I might actually learn how it is done. Thank you in advance! 1) Solve. Check for extraneous solutions √3x+1-√x+1=2 1)2,6 2) 0,8 3) 1 4) 8

Directrix (directrix):

Is this the problem: √(3x+1) - √(x+1) = 2

OpenStudy (bluepatriot):

Yes, my dear. An answer is appreciated but a step by step explanation would really save my life. I have a few more like these and I need another explanation on how to do them.

OpenStudy (bluepatriot):

@Directrix

OpenStudy (jiteshmeghwal9):

\[(\sqrt{3x+1}-\sqrt{x+1}=2)\]square both sides\[(3x+1)+(x+1)-2(\sqrt{3x+1})(\sqrt{x+1})=4\]\[2\sqrt{3x^2+4x+1}=4x-2\]gt it upto this step ?

OpenStudy (jiteshmeghwal9):

next step is to square it again. finally u will gt a quadratic equation. then u can solve it using quadratic formula, factorization or perfect square method.

OpenStudy (bluepatriot):

I think I do.